OpenStudy (twizttiez):

Use the graph to state the solution for the system. x – y = 4 (line a) x + y = 4 (line b) A. (0, 4) B. (4, 0) C. (0, –4) D. (–4, 0)

OpenStudy (mathmath333):

what is the point where they intesect ?

OpenStudy (twizttiez):

0,4 right?

OpenStudy (mathmath333):

\(\Huge~no!\)

OpenStudy (twizttiez):

Ah! ok 4,0?

OpenStudy (mathmath333):

\(\small yes\)

OpenStudy (twizttiez):

Can you help with a few more?

OpenStudy (mathmath333):

ya bring it on

OpenStudy (twizttiez):

Yay!

OpenStudy (twizttiez):

Do you want me to open a new question?

OpenStudy (mathmath333):

i dont care

OpenStudy (twizttiez):

How many solutions are there to the following system of equations? 3x – 9y = 0 –x + 3y = –3 A. 1 B. 2 C. 0 D. infinitely many

OpenStudy (twizttiez):

C.0?

OpenStudy (twizttiez):

???

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{3x – 9y = 0--\color{red}{(1)}\\~\\ –x + 3y = –3---\color{red}{(2)}}\end{align}\) divide (1) by 3 then add with (2)

OpenStudy (twizttiez):

That would make it inconsistent right?

OpenStudy (mathmath333):

no

OpenStudy (twizttiez):

Oh ok then hmm

OpenStudy (twizttiez):

Would it be D?

OpenStudy (mathmath333):

first add and find x and y

OpenStudy (twizttiez):

I'm confused

OpenStudy (twizttiez):

2x-y = 5; 3x+2y = 4; --> x= 2; y=-1; 3c+5a = 115; c+a = 33; --> c= 25; a= 8; 3c+5a = 150; c+a = 40; --> c= 25; a= 15; 6x-9y = 9; 2x-3y = 6; 1 -3/2 -3/2 0 0 -3

OpenStudy (twizttiez):

Right?

OpenStudy (twizttiez):

So therefore it would be inconsistent

OpenStudy (twizttiez):

I think

OpenStudy (mathmath333):

do u think ur work and ur question matches

OpenStudy (twizttiez):

Oh o_O ok one moment

OpenStudy (twizttiez):

3x-9y = 0; -x+3y = -3; 1 -3 0 0 0 3 ?

OpenStudy (twizttiez):

Ok well i will brb

OpenStudy (twizttiez):

Ok im back

OpenStudy (mathmath333):

well the system of equtions of the form \(ax+by=0\\~\\ mx+ny=c\) where and \(c\neq 0\) and either \(a\neq m\) or \(b\neq n\) has \(2\) solutions

OpenStudy (twizttiez):

Ohh ok thanks i do have like 8 more wanna help?